# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Bulk-Surface Interaction

### v2

@@ 4  Conclusion
Starting from local mass and energy balance laws, we systematically derived the governing equations and boundary conditions using the Lagrange multiplier method. The resulting system satisfies the fundamental physical principles of mass conservation, local energy dissipation, and preservation of the bulk volume and surface area of the evolving domain. We further showed that several classical and recently proposed models (cf.[5, 50]) appear as special limiting cases of our formulation. This highlights the generality and flexibility of the present model.
@@ Acknowledgement
Both authors were partially supported by the “RTG 2339” IntComSin funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 321821685. PK was additionally supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 52469428. YL was partially supported by the startup funding from Nanjing Normal University, the Natural Science Foundation of Jiangsu Province (Grant No. BK20240572), the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (Grant No. 24KJB110020), and the China Postdoctoral Science Foundation under Grant Number 2025M773078. The support is gratefully acknowledged.

### v3

@@ 4  Conclusion
Starting from local mass balance laws, we systematically derived the governing equations and boundary conditions in two different ways: by the Lagrange Multiplier Approach and, assuming matched densities and no mass flux between bulk and surface, by the Energetic Variational Approach. The resulting system satisfies the fundamental physical principles of mass conservation, local energy dissipation, and preservation of the bulk volume and surface area of the evolving domain. We further showed that several classical and recently proposed models (cf.[Abels2012, Giorgini2023]) appear as special limiting cases of our formulation. This highlights the generality and flexibility of the present model.
@@ Acknowledgement
We thank the anonymous referees for their careful reading and valuable suggestions, which helped us to improve the manuscript. Both authors were partially supported by the “RTG 2339” IntComSin funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 321821685. PK was additionally supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 52469428. YL was partially supported by the startup funding from Nanjing Normal University, the Natural Science Foundation of Jiangsu Province (Grant No. BK20240572), the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (Grant No. 24KJB110020), and the China Postdoctoral Science Foundation under Grant Number 2025M773078. The support is gratefully acknowledged.